# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v2

@@ I Introduction
Mechanically interlocked polymers (MIPs) are a class of polymer architectures distinguished by components linked with mechanical entanglements rather than covalent bonds. The schematic representation in Fig. 1 shows different classes of MIPs in comparison with a typical linear polymer (Fig. 1a). Polyrotaxanes (Fig. 1b) consist of multiple ring-shaped molecules threaded onto a linear backbone, held in place by bulky end caps called stoppersHarrison and Harrison (1967); Wenz and Keller (1992). Daisy chains (Fig. 1c) are made from covalently-bonded rings and linear sections interlocked in seriesAshton et al. (1998); Wenz et al. (2006). Polycatenanes (Fig. 1d) are polymers made of interlocked rings, like a chainSchill and Lüttringhaus (1964); Frisch and Wasserman (1961); Weidmann et al. (1999); Wu et al. (2017). These mechanical bonds impart MIPs with increased conformational freedom, setting them apart from typical polymers Hart et al. (2021) and leading to unique propertiesSluysmans and Stoddart (2019); Liu et al. (2022) such as shape memory, ductile glassy states, large strain-at-break, scratch resistance, and stimuli-responsive mechanical properties Hart et al. (2021); Zhou et al. (2024); Ando and Ito (2025). From theoretical studies it is anticipated that MIPs will show large loss modulus, low activation energy for viscous flow, and rapid stress relaxation Weidmann et al. (1999), potentially acting as superior energy-damping materials with excellent toughness Wu et al. (2017). In addition, equilibrium and strain simulations of rotaxanes and rotaxane-based materials, like slide-ring gels and tendomer gels, have demonstrated exceptional elasticity and yield stresses surpassing conventional cross-linked polymer gels Li et al. (2021); Yasuda et al. (2019); Chen et al. (2023); Tanahashi and Koga (2021); Uehara et al. (2022); Yasuda et al. (2020); Müller et al. (2019, 2021, 2022). These examples demonstrate the unique properties imparted by mechanical bonds.
@@
Although mechanically interlinking molecules is a synthetic challenge Segawa et al. (2019); Leigh et al. (2014), recent advances have enabled the production of MIPs with moderate molecular weights. The largest polyrotaxane synthesized to date has weight-average molecular weight $M_{w}$ = 119,000 g/mol with 86 threaded rings Araki et al. (2005) while the largest daisy chain polymer has 11 repeat units and number-average molecular weight $M_{n}\approx$ 25,000 g/mol Cai et al. (2020). Ultra-high-molecular-weight polycatenanes have recently been synthesized, consisting of up to 2,800 rings with a molecular weight of 4.5 MDa Wang et al. (2025). Nano-scale polycatenanes comprising up to 22 interlocked rings have been synthesized through the self-assembly of supramolecular toroidal building blocks Datta et al. (2020).
@@
Despite advances in synthesis, the relationship between mechanical bonds and flow properties is largely unexplored. Computational techniques present an effective method to link the molecular structure, dynamics, and macroscopic behavior of polymersGartner III and Jayaraman (2019); Li et al. (2013); Gedde et al. (2021); Seyedi et al. (2020); Kremer and Grest (1990). In the realm of MIPs, simulations have quantified the structural and dynamic properties of doubly interlocked rings (2-catenanes) in equilibrium conditions Rauscher et al. (2020); Bohn et al. (2010), in shear Li et al. (2025), and in glassy states Li et al. (2023). Simulation studies have explored the equilibrium behavior of longer polycatenanes (up to 300 rings), isolated Tubiana et al. (2022); Chiarantoni and Micheletti (2022); Rauscher et al. (2018); Dehaghani et al. (2020); Pakula and Jeszka (1999); Li et al. (2021) and in a melt state Rauscher et al. (2020, 2020); Hagita et al. (2021), uncovering characteristics that blend features of ring and linear polymers. In addition, computational investigations have probed the structure and dynamics of catenanes under confinement Chiarantoni and Micheletti (2023) and during pore translocation Caraglio et al. (2017); Suma and Micheletti (2017) and rotaxanes have been analyzed in equilibrium Li et al. (2021); Yasuda et al. (2019). These simulations have highlighted the unique behavior of MIPs compared to conventional polymers, demonstrating how structural features, such as ring size and the number of mechanical bonds, influence properties like chain stiffness Pakula and Jeszka (1999); Rauscher et al. (2020), $\theta$-temperature Dehaghani et al. (2020); Guo et al. (2023), and relaxation dynamics Pakula and Jeszka (1999); Rauscher et al. (2018). Importantly, these studies show that MIPs exhibit emergent properties not observed in typical linear and cyclic polymer architecturesPakula and Jeszka (1999), highlighting the crucial role of topology in governing their behavior.
@@
Here we simulate dilute solutions of MIPs with coarse-grained Brownian dynamics, an approach that has been effective at characterizing the rheology of many polymer morphologies, like linear Lyulin et al. (1999), combs Mai et al. (2018), bottlebrushes Dutta and Sing (2024), and ring polymers Young et al. (2019). We report rheological properties for polyrotaxanes, daisy chains, and polycatenanes in both steady shear and steady extensional flow, a critically uncharacterized domain for MIPs. We show that all three MIP types show enhanced tumbling in shear flow, weaker shear thinning, and lower normal stress differences in extensional flow. Also, while polyrotaxanes show higher viscosity both in shear and extension, polycatenanes and daisy chains exhibit lower viscosities. Polymer extension is typically greater in both shear and extensional flows, and approaches linear values as the density of mechanical bonds in MIPs decreases.
@@ II Methods
We simulated single polymers in the dilute solution limit with a coarse-grained bead-based model. Using a previously established model for MIPs Rauscher et al. (2020); Becerra et al. (2024), each polymer consists of $N$ beads that interact through bonded and non-bonded potentials. The pairwise non-bonded interaction potential is the Weeks–Chandler–Anderson potential Weeks et al. (1971)
@@
where $r_{ij}=\left|\mathbf{r}_{j}-\mathbf{r}_{i}\right|$ is the distance between particles $i$ and $j$, $r_{c}=2.2449241$ is the minimum energy cutoff distance, $U_{c}=4\epsilon\left[\left(\frac{\sigma}{r_{c}}\right)^{12}-\left(\frac{\sigma}{r_{c}}\right)^{6}\right]$ is the Lennard-Jones energy at the cutoff distance, and $\sigma$ and $\epsilon$ are the Lennard-Jones length and energy parameters, respectively. To prevent polyrotaxane and daisy chain rings from dethreading, we introduced non-bonded interactions between capping beads and ring beads that are modeled without the attractive term, shift and cutoff by using $U_{\mathrm{cap}}=4\epsilon_{\mathrm{cap}}\left(\frac{\sigma}{r_{ij}}\right)^{12}$, where $\epsilon_{\mathrm{cap}}=10^{5}\epsilon$ in daisy chains and $\epsilon_{\mathrm{cap}}=10^{10}\epsilon$ in polyrotaxanes. Due to the absence of an attractive term, these $\epsilon_{\mathrm{cap}}$ values are equivalent to 2.6 and 6.8 fold increase in $\sigma$ size parameter, respectively. Bonds between beads are modeled with finitely extensible nonlinear elastic (FENE) springs Kremer and Grest (1990):
@@
The bead positions evolve according to the following stochastic differential equation incorporating hydrodynamic interactions and implicit solvent Ermak and McCammon (1978); Ottinger (1996):
@@
is the Rotne–Prager–Yamakawa tensor Rotne and Prager (1969); Yamakawa (1970). To remove the overall translational motion of the molecule, the solvent velocity is centered around the center of mass of the polymer $\mathbf{r}_{\mathrm{COM}}=[x_{\mathrm{COM}},y_{\mathrm{COM}},z_{\mathrm{COM}}]^{\top}$ Miao et al. (2017). The solvent velocity $\mathbf{v}_{i}$ for bead $i$ with spatial coordinate $\mathbf{r}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ is zero at equilibrium; $\dot{\gamma}(y_{i}-y_{\mathrm{COM}}),0,0]^{\top}$ in shear flow, where $\dot{\gamma}$ is the shear rate; and $\dot{\varepsilon}(x_{i}-x_{\mathrm{COM}}),-\dot{\varepsilon}(y_{i}-y_{\mathrm{COM}})/2,-\dot{\varepsilon}(z_{i}-z_{\mathrm{COM}})/2]^{\top}$ in uniaxial extensional flow, where $\dot{\varepsilon}$ is the extension rate. Furthermore, the Weissenberg number (Wi) measures the flow strength and is defined as $\dot{\gamma}\tau_{R}$ for shear flow and $\dot{\varepsilon}\tau_{R}$ for uniaxial extensional flow, where $\tau_{R}$ is the longest relaxation time of the polymer. Relaxation time is $\tau_{R}=R_{g}^{2}/6D_{\mathrm{COM}}$, where the center-of-mass diffusivity was calculated from the equilibrium conformations using the Kirkwood formalism Dünweg et al. (2002). Relaxation time results are summarized in SI Tables A.1 to A.4. The variables are nondimensionalized using $a$ as the unit of length, $k_{B}T$ as the unit of energy, and the diffusion time of a bead $\tau=\zeta a^{2}/{k_{B}T}$ as the unit of time. Eq. 3 is integrated numerically using a semi-implicit Euler scheme Jendrejack et al. (2000, 2002) to allow the use of larger time steps compared to explicit methods. The diffusivity matrix $\mathbf{D}$ and the Brownian term are treated explicitly and the remaining terms are treated implicitly. The following set of nonlinear algebraic equations in $\mathbf{r}^{*}$ are obtained using the semi-implicit algorithm:
@@
Equation 6 is solved iteratively by a Jacobian-free Newton–Krylov method Knoll and Keyes (2004) through the NITSOL software package Pernice and Walker (1998) without using a preconditioner. This approach has recently been used to simulate bottlebrush polymers Dutta and Sing (2024). We use OpenMP shared memory parallelism with 1-4 CPUs per simulation (depending on system size) to more efficiently calculate forces, perform the Cholesky decomposition of the $\mathbf{D^{*}}$ matrix, and integrate the equation of motion. Initial configurations were generated with equilibrium bond lengths and energy minimization was carried out prior to production simulations. All properties were calculated by averaging over 10 to 20 trajectories of at least 100 relaxation times long ($10^{6}-10^{8}$ time steps) with time steps ranging from $10^{-2}$ to $10^{-4}$ depending on the strain rate. Error bars in the graphs represent the standard error of the mean. We calculated power-law exponents and their error using parametric bootstrap (Monte Carlo resampling) Efron (1992). The fractional extension in the flow direction is $\langle X_{bb}\rangle/L$, where $X_{bb}$ is the difference between minimum and maximum coordinates in the $x$ direction and $L$ is the maximum contour length. The maximum contour length is L=3a(N1)L=3a(N-1) and $L=3a(N_{bb}-1)$ for linear polymers and polyrotaxanes respectively, where $N$ is the total number of beads and $N_{bb}$ is the number of backbone beads. For daisy chains and polycatenanes, a pair of equal and opposite forces is applied to the end beads of the polymers in the $x$-direction and gradually increased under zero-temperature conditions. The polymer extension just before a mechanical bond fails, corresponding to a maximum in stress ($\tau^{p}_{xx}$) is recorded as the maximum contour length, $L$.
In dilute polymer solutions, the polymer contribution to the stress tensor ($\bm{\tau}^{p}$) is calculated using the following equation Bird et al. (1977); Graham (2018); Bosko and Ravi Prakash (2008)
@@
Under steady uniaxial extension, we calculate the fractional extension $\langle X_{bb}\rangle/L$, extensional viscosity $\eta_{e}$ and first normal stress difference $N_{1}$ across a range of Wi values. Under steady shear flow, we calculate the fractional extension in $x$ (flow), $y$ (gradient), and $z$ (cross-flow) directions ($\langle X_{bb}\rangle/L$, $\langle Y_{bb}\rangle/L$, $\langle Z_{bb}\rangle/L$), first normal stress coefficient ($\psi_{1}$), shear viscosity $\eta$, orientation angle with respect to the flow $\theta$, and tumbling frequency ($\omega\tau_{R}$) over a range of Wi values. In shear, we track and analyze the evolution of the angle between the end-to-end vector of the polymer chains and the $x$-axis to calculate the tumbling time, similar to the work of Dalal et. al. Saha Dalal et al. (2012). We calculate tumbling frequency as inverse tumbling time, normalized by the relaxation time ($\tau_{R}$).
@@ III Results
Figure 2: Polyrotaxanes properties in extensional flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) first normal stress difference $(N_{1})$, (c) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
Figure 3: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.

### v3

@@ I Introduction
Mechanically interlocked polymers (MIPs) are a class of polymer architectures distinguished by components linked with mechanical entanglements rather than covalent bonds. The schematic representation in Fig. 1 shows different classes of MIPs in comparison with a typical linear polymer (Fig. 1a). Polyrotaxanes (Fig. 1b) consist of multiple ring-shaped molecules threaded onto a linear backbone, held in place by bulky end caps called stoppersharrison1967synthesis; wenz1992threading. Daisy chains (Fig. 1c) are made from covalently-bonded rings and linear sections interlocked in seriesashton1998supramolecular; wenz2006cyclodextrin. Polycatenanes (Fig. 1d) are polymers made of interlocked rings, like a chainschill1964preparation; frisch1961chemical; weidmann1999poly; wu2017poly. These mechanical bonds impart MIPs with increased conformational freedom, setting them apart from typical polymers hart2021material and leading to unique propertiessluysmans2019burgeoning; liu2022polycatenanes such as shape memory, ductile glassy states, large strain-at-break, scratch resistance, and stimuli-responsive mechanical properties hart2021material; zhou2024mechanically; ando2025recent. From theoretical studies it is anticipated that MIPs will show large loss modulus, low activation energy for viscous flow, and rapid stress relaxation weidmann1999poly, potentially acting as superior energy-damping materials with excellent toughness wu2017poly. In addition, equilibrium and strain simulations of rotaxanes and rotaxane-based materials, like slide-ring gels and tendomer gels, have demonstrated exceptional elasticity and yield stresses surpassing conventional cross-linked polymer gels li2021sliding; yasuda2019sliding; chen2023elasticity; tanahashi2021molecular; uehara2022molecular; yasuda2020molecular; muller2019tendomers; muller2021swelling; muller2022elasticity. These examples demonstrate the unique properties imparted by mechanical bonds.
@@
Although mechanically interlinking molecules is a synthetic challenge segawa2019topological; leigh2014star, recent advances have enabled the production of MIPs with moderate molecular weights. The largest polyrotaxane synthesized to date has weight-average molecular weight $M_{w}$ = 119,000 g/mol with 86 threaded rings araki2005efficient while the largest daisy chain polymer has 11 repeat units and number-average molecular weight $M_{n}\approx$ 25,000 g/mol cai2020molecular. Ultra-high-molecular-weight polycatenanes have recently been synthesized, consisting of up to 2,800 rings with a molecular weight of 4.5 MDa wang2025fully. Nano-scale polycatenanes comprising up to 22 interlocked rings have been synthesized through the self-assembly of supramolecular toroidal building blocks datta2020self.
@@
Despite advances in synthesis, the relationship between mechanical bonds and flow properties is largely unexplored. Computational techniques present an effective method to link the molecular structure, dynamics, and macroscopic behavior of polymersgartner2019modeling; li2013challenges; Gedde2021; seyedi2020initiator; kremer1990dynamics. In the realm of MIPs, simulations have quantified the structural and dynamic properties of doubly interlocked rings (2-catenanes) in equilibrium conditions rauscher2020hydrodynamic; bohn2010influence, in shear li2025effects, and in glassy states li2023glass. Simulation studies have explored the equilibrium behavior of longer polycatenanes (up to 300 rings), isolated tubiana2022circular; chiarantoni2022effect; rauscher2018topological; dehaghani2020effects; pakula1999simulation; li2021double and in a melt state rauscher2020thermodynamics; rauscher2020dynamics; hagita2021mathematical, uncovering characteristics that blend features of ring and linear polymers. In addition, computational investigations have characterized the linear and nonlinear elasticity of single polycatenanes chen2023topological; chen2024nonlinear and probed the structure and dynamics of catenanes under confinement chiarantoni2023linear and during pore translocation caraglio2017driven; suma2017pore. Simulations have also been used to analyze rotaxanes in equilibrium li2021sliding; yasuda2019sliding and daisy chains formed in ring polymer melts under extensional flow o2020topological. These simulations have highlighted the unique behavior of MIPs compared to conventional polymers, demonstrating how structural features, such as ring size and the number of mechanical bonds, influence properties like chain stiffness pakula1999simulation; rauscher2020thermodynamics, elasticity chen2023topological; chen2024nonlinear, $\theta$-temperature dehaghani2020effects; guo2023theta, and relaxation dynamics pakula1999simulation; rauscher2018topological. Importantly, these studies show that MIPs exhibit emergent properties not observed in typical linear and cyclic polymer architecturespakula1999simulation, highlighting the crucial role of topology in governing their behavior.
@@
Here we simulate dilute solutions of MIPs with coarse-grained Brownian dynamics, an approach that has been effective at characterizing the rheology of many polymer morphologies, like linear lyulin1999brownian, combs mai2018stretching, bottlebrushes dutta2024brownian, and ring polymers young2019ring. We report rheological properties for polyrotaxanes, daisy chains, and polycatenanes in both steady shear and steady extensional flow, a critically uncharacterized domain for MIPs. We show that all three MIP types show enhanced tumbling in shear flow, weaker shear thinning, and lower normal stress differences in extensional flow. Also, while polyrotaxanes show higher viscosity both in shear and extension, polycatenanes and daisy chains exhibit lower viscosities. Polymer extension is typically greater in both shear and extensional flows, and approaches linear values as the density of mechanical bonds in MIPs decreases.
@@ II Methods
We simulated single polymers in the dilute solution limit with a coarse-grained bead-based model. Using a previously established model for MIPs rauscher2020thermodynamics; becerra2024single, each polymer consists of $N$ beads that interact through bonded and non-bonded potentials. The pairwise non-bonded interaction potential is the Weeks–Chandler–Anderson potential weeks1971role
@@
where $r_{ij}=\left|\mathbf{r}_{j}-\mathbf{r}_{i}\right|$ is the distance between particles $i$ and $j$, $r_{c}=2.2449241$ is the minimum energy cutoff distance, $U_{c}=4\epsilon\left[\left(\frac{\sigma}{r_{c}}\right)^{12}-\left(\frac{\sigma}{r_{c}}\right)^{6}\right]$ is the Lennard-Jones energy at the cutoff distance, and $\sigma$ and $\epsilon$ are the Lennard-Jones length and energy parameters, respectively. To prevent polyrotaxane and daisy chain rings from dethreading, we introduced non-bonded interactions between capping beads and ring beads that are modeled without the attractive term, shift and cutoff by using $U_{\mathrm{cap}}=4\epsilon_{\mathrm{cap}}\left(\frac{\sigma}{r_{ij}}\right)^{12}$, where $\epsilon_{\mathrm{cap}}=10^{5}\epsilon$ in daisy chains and $\epsilon_{\mathrm{cap}}=10^{10}\epsilon$ in polyrotaxanes. Due to the absence of an attractive term, these $\epsilon_{\mathrm{cap}}$ values are equivalent to 2.6 and 6.8 fold increase in $\sigma$ size parameter, respectively. Bonds between beads are modeled with finitely extensible nonlinear elastic (FENE) springs kremer1990dynamics:
@@
The bead positions evolve according to the following stochastic differential equation incorporating hydrodynamic interactions and implicit solvent ermak1978brownian; ottinger1996stochastic:
@@
is the Rotne–Prager–Yamakawa tensor rotne1969variational; yamakawa1970transport. To remove the overall translational motion of the molecule, the solvent velocity is centered around the center of mass of the polymer $\mathbf{r}_{\mathrm{COM}}=[x_{\mathrm{COM}},y_{\mathrm{COM}},z_{\mathrm{COM}}]^{\top}$ miao2017iterative. The solvent velocity $\mathbf{v}_{i}$ for bead $i$ with spatial coordinate $\mathbf{r}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ is zero at equilibrium; $\dot{\gamma}(y_{i}-y_{\mathrm{COM}}),0,0]^{\top}$ in shear flow, where $\dot{\gamma}$ is the shear rate; and $\dot{\varepsilon}(x_{i}-x_{\mathrm{COM}}),-\dot{\varepsilon}(y_{i}-y_{\mathrm{COM}})/2,-\dot{\varepsilon}(z_{i}-z_{\mathrm{COM}})/2]^{\top}$ in uniaxial extensional flow, where $\dot{\varepsilon}$ is the extension rate. Furthermore, the Weissenberg number (Wi) measures the flow strength and is defined as $\dot{\gamma}\tau_{R}$ for shear flow and $\dot{\varepsilon}\tau_{R}$ for uniaxial extensional flow, where $\tau_{R}$ is the longest relaxation time of the polymer. Relaxation time is $\tau_{R}=R_{g}^{2}/6D_{\mathrm{COM}}$, where the center-of-mass diffusivity was calculated from the equilibrium conformations using the Kirkwood formalism dunweg2002corrections. Relaxation time results are summarized in SI Tables A.1 to A.4. The variables are nondimensionalized using $a$ as the unit of length, $k_{B}T$ as the unit of energy, and the diffusion time of a bead $\tau=\zeta a^{2}/{k_{B}T}$ as the unit of time. Eq. 3 is integrated numerically using a semi-implicit Euler scheme jendrejack2000hydrodynamic; jendrejack2002stochastic to allow the use of larger time steps compared to explicit methods. The diffusivity matrix $\mathbf{D}$ and the Brownian term are treated explicitly and the remaining terms are treated implicitly. The following set of nonlinear algebraic equations in $\mathbf{r}^{*}$ are obtained using the semi-implicit algorithm:
@@
Equation 6 is solved iteratively by a Jacobian-free Newton–Krylov method knoll2004jacobian through the NITSOL software package pernice1998nitsol without using a preconditioner. This approach has recently been used to simulate bottlebrush polymers dutta2024brownian. We use OpenMP shared memory parallelism with 1-4 CPUs per simulation (depending on system size) to more efficiently calculate forces, perform the Cholesky decomposition of the $\mathbf{D^{*}}$ matrix, and integrate the equation of motion. Initial configurations were generated with equilibrium bond lengths and energy minimization was carried out prior to production simulations. All properties were calculated by averaging over 10 to 20 trajectories of at least 100 relaxation times long ($10^{6}-10^{8}$ time steps) with time steps ranging from $10^{-2}$ to $10^{-4}$ depending on the strain rate. Error bars in the graphs represent the standard error of the mean. We calculated power-law exponents and their error using parametric bootstrap (Monte Carlo resampling) efron1992bootstrap. The fractional extension in the flow direction is $\langle X\rangle/L$, where $X$ is the difference between minimum and maximum coordinates in the $x$ direction and $L$ is the maximum contour length. The maximum contour length is L=3a(N1)L=3a(N-1) and $L=3a(N_{bb}-1)$ for linear polymers and polyrotaxanes respectively, where $N$ is the total number of beads and $N_{bb}$ is the number of backbone beads. For daisy chains and polycatenanes, a pair of equal and opposite forces is applied to the end beads of the polymers in the $x$-direction and gradually increased under zero-temperature conditions. The polymer extension just before a mechanical bond fails, corresponding to a maximum in stress ($\tau^{p}_{xx}$) is recorded as the maximum contour length, $L$.
In dilute polymer solutions, the polymer contribution to the stress tensor ($\bm{\tau}^{p}$) is calculated using the following equation bird1977dynamics; Graham_2018; bosko2008effect
@@
Under steady uniaxial extension, we calculate the fractional extension $\langle X\rangle/L$, extensional viscosity $\eta_{e}$ and first normal stress difference $N_{1}$ across a range of Wi values. Under steady shear flow, we calculate the fractional extension in $x$ (flow), $y$ (gradient), and $z$ (vorticity) directions ($\langle X\rangle/L$, $\langle Y\rangle/L$, $\langle Z\rangle/L$), first normal stress coefficient ($\psi_{1}$), shear viscosity $\eta$, orientation angle with respect to the flow $\theta$, and tumbling frequency ($\omega\tau_{R}$) over a range of Wi values. In shear, we track and analyze the evolution of the angle between the end-to-end vector of the polymer chains and the $x$-axis to calculate the tumbling time, similar to the work of Dalal et. al. saha2012tumbling. We calculate tumbling frequency as inverse tumbling time, normalized by the relaxation time ($\tau_{R}$).
We note that the parameters of this model allow the mechanical bonds to maintain integrity up to moderate flow strength in extensional flow and high flow strength in shear flow. For strong enough flows, however, the hydrodynamic forces cause bond crossing events that break up the mechanical interlocks. We report data for flow strengths up until the point that the mechanical bonds fail. Mechanical bond failure in MIPs is expected from experiments muramatsu2021rotaxane; zhang2019mechanical; lee2016mechanical, and we address this point in greater detail in our discussion.
@@ III Results
Figure 2: Polyrotaxanes properties in extensional flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) first normal stress difference $(N_{1})$, (c) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
Figure 3: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.